OpenStudy (anonymous):

I have some math questions that I'm having a little trouble figuring out. If someone could please help me, I will medal and all help is appreciated! I have to write an equation for each condition listed. 1. The line with a slope of 7 and a y-intercept of -1 (in slope-intercept form) 2. The line that passes through (1, -5) and has a slope of 3 ( in point-slope form) 3. The vertical line that passes through (-7, 4) 4. The horizontal line that passes through (-7, 4) Thanks again!

OpenStudy (mathmale):

As a start, would you please type in the "slope-intercept form" of the equation of a straight line?

OpenStudy (anonymous):

I'm sorry, but what do you mean?

OpenStudy (welshfella):

Have yo seen this before? y = mx + b

OpenStudy (anonymous):

I think so

OpenStudy (welshfella):

Its the slope intercept form of a straight line do you know what m and b are?

OpenStudy (anonymous):

Oh okay and no I don't I really don't understand math :/

OpenStudy (welshfella):

This is what is caalled a 'general form'. m is the slope of the line and b is the y-intercept y = b when x = 0:- y = m(0) + b = 0 + b = b

OpenStudy (anonymous):

Okay

OpenStudy (welshfella):

examples of this type are y = 2x + 3 y = 0.5x - 3

OpenStudy (welshfella):

If you look at your number one you can find the equation by replacing m by the given slope and b by the given y-intercept.

OpenStudy (anonymous):

Ok...

OpenStudy (welshfella):

m = 7 and b = -1 so y = ?

OpenStudy (123ab456c):

@Speedy2197 YOU LOOK LIKE ARIEL! FROM THE MOVIE THE LITTLE MERMAID

OpenStudy (anonymous):

Y would = I'm not sure... Like I understand what you're saying, but I'm just not sure what it would =

OpenStudy (anonymous):

@123AB456C Thanks lol My real name is actually Ayriel Ariel, but with a y

OpenStudy (welshfella):

y = mx + b just replace m by 7 and b by -1

OpenStudy (anonymous):

Okay and what does the x represent in the mx?

OpenStudy (welshfella):

the x and y are variable y depends on the value of x and vice versa just worry about the costants in the equations - m and b

OpenStudy (anonymous):

Okay

OpenStudy (welshfella):

all you have to do is replace m and b in the general equation by 7 and -1... its easy

OpenStudy (anonymous):

I understand that problem now, thank you. You explained it well

OpenStudy (welshfella):

yw

OpenStudy (anonymous):

How do I solve number 2,3, and 4 though?

OpenStudy (welshfella):

first what is the answer to number 1?

OpenStudy (anonymous):

m=7 b = -1 y=mx+b

OpenStudy (welshfella):

but you havent replaced the m and b in the equation

OpenStudy (welshfella):